# Efficient Exploration and Discriminative World Model Learning with an Object-Centric Abstraction

### v1

@@ Abstract
We demonstrate the model’s ability to (i) efficiently solve single tasks, (ii) transfer zero-shot and few-shot across item types and environments, and (iii) plan across long horizons. Across a suite of 2D crafting and MiniHack environments, we empirically show our model significantly out-performs state-of-the-art low-level methods (without abstraction), as well as performant model-free and model-based methods using the same abstraction. Finally, we show how to reinforce learn low level object-perturbing policies, as well as supervise learn the object mapping itself.
@@ 1 Introduction
Since the inception of reinforcement learning (RL), the problems of exploration and world model learning have been major roadblocks. RL requires an agent to learn both basic motor abilities, and explore long sequences of interactions in the environment. Currently, we have made tremendous progress on problems of low-level control for short-horizon problems. With well defined single tasks, given enough samples (or demonstrations), there are well-developed ways of training low level policies reliably (Schulman et al.,, 2017; Hafner et al.,, 2023).
Perhaps a better way to explore complex environments is to treat the high-level exploration problem separately. We increasingly have the ability to get semantically rich abstractions: via representation learning (Locatello et al.,, 2020), object segmentation (Kirillov et al.,, 2023), and visual-language models  (Alayrac et al.,, 2022); even in difficult control settings such as robotics (Liu et al.,, 2024). In addition, with the recent explosion in the field of natural language processing, people are increasingly interested in addressing RL environments on a semantic level (Andreas et al.,, 2017; Jiang et al.,, 2019; Chen et al.,, 2021). Thus, we ask a timely question: if the agent has a good ability to perceive objects, what more can we do other than tabula rasa learning at the level of pixels and primitive motor actions?
@@
The main results of this work investigates decision making at the semantic level as defined by the Ab-MDP, assuming access to an object centric map and object perturbing policies. We empirically evaluate our method in a set of 2D crafting games, and MiniHack environments that are known to be very difficult to explore (Henaff et al.,, 2022). We then show how to learn components of the Ab-MDP when it is not given: by reinforcement learning object perturbation policies given an object mapping, and supervised learning of the object centric mapping itself.
@@ 2 Problem Setting
We begin by considering a (reward free) “base” Markov Decision Process (MDP) as the tuple $\langle\mathcal{S},\mathcal{A},P\rangle$, with state space $\mathcal{S}$, primitive action space $\mathcal{A}$, and transition probability function $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ (Puterman,, 1994). Below we construct the “abstracted” space.
@@ 2.2 Abstracted Item-Attribute MDP
Specifically, the Ab-MDP is the tuple $\langle\mathcal{X},\mathcal{B},\textbf{T}\rangle$, with abstract states $\mathcal{X}$, abstract behaviours $\mathcal{B}$, and transition probability function $\textbf{T}:\mathcal{X}\times\mathcal{B}\times\mathcal{X}\rightarrow[0,1]$. $\textbf{T}(X^{\prime}|X,b)$ describes the probability of being in abstract state $X^{\prime}$ when starting from $X$, executing behaviour $b$, and waiting until an abstract transition occurs. Figure 1 provides an example of an abstract transition and associated low level steps. The Ab-MDP can be interpreted as Options (Sutton et al.,, 1999; Precup,, 2000), which we discuss in Appendix B.1.1.
@@ 3 Methods
While any standard RL methods can be used to solve the Ab-MDP, we develop a method to more fully make use of the item-attribute structure and item-perturbing behaviours. Specifically, we design a model-based method that efficiently explores and learn an accurate full world model. Our method, Model-based Exploration of abstracted Attribute Dynamics (MEAD) can be broken down into three components: (i) A forward model $f_{\theta}(X,b)$, (ii) A reward function $r(X,b)$, (iii) A planner which maximizes the reward function.
@@ 3.1 Forward model
A forward model takes in a current abstract state and behaviour to model the distribution over next states, $X_{t+1}\sim p(X_{t+1}|X_{t},b_{t})$. Typically, this distribution has support over the full state space $\mathcal{X}$. We observe that by construction, Ab-MDP considers competent policies as ones that change single item-attributes. Thus, we make the modelling assumption that an abstract transition would result in either (i) a single object’s attribute changing (as specified by the executed behaviour), or (ii) the abstract state remaining unchanged. It is of course possible for multiple items’ attributes to change at once, or for an unexpected item’s attribute to change—we show with a competent set of policies and re-planning our model is robust to this in practice in Appendix B.2.2.
@@
We use model function class $f_{\theta}:\mathcal{X}\times\mathcal{B}\rightarrow[0,1]$ with trainbale parameters $\theta$, and fit it to approximate success probability (Equation 1). We can estimate the empirical success probability from a dataset of trajectories simply by counting, and fit a binary cross entropy loss. We found this to be a stable, simple-to-optimize objective. More details in Appendix C.1.
@@ 3.2 Planning for exploration
(b) Dijkstra planning at evaluation time to maximize success probability of reaching any goal state
@@
We want the agent to explore indefinitely to discover how the world works and represent it in its model $f_{\theta}$. We use a count-based intrinsic reward introduced in Zhang et al., (2018) which encourages the agent to uniformly cover all state-behaviours (details in Appendix C.3):
@@ 4 Results
We use two sets of of environment. One is 2D crafting, which requires agent to craft objects following a Minecraft-like crafting tree with each crafting stage requiring multiple low level actions (Chen et al.,, 2020). Second is MiniHack (Samvelyan et al.,, 2021), which contains a series of extremely difficult exploration games (Henaff et al.,, 2022). We detail the environments further in Appendix D.
For baselines, we use strong model-free (asynch PPO, (Schulman et al.,, 2017; Petrenko et al.,, 2020)) and model-based (Dreamer-v3, (Hafner et al.,, 2023; Sekar et al.,, 2020)) baselines. More details in Appendix E. All the methods are trained in the Ab-MDP. We report low level steps and plot 95 confidence interval. See Appendix G for other experimental details.
@@ 4.1 Learning from scratch in single environments
Figure 4: Results in games trained from scratch in Ab-MDP. Triangles and stars denote low-level only methods. For MiniHack, we also show the final performance of state-of-the-art exploration methods in the low level MDP (star/triangle), from Henaff et al., (2022). X-axis on log scale.
Figure 4 shows the from-scratch Ab-MDP training results for strong baselines and our method. We see that Dreamer-v3, a state-of-the-art model-based method, is highly sample efficient. PPO, a model-free method, stably optimizes the reward, but at a data budget of more than two orders of magnitude (hundreds of thousands versus 10 million). Our method learns the Ab-MDP with an additional order of magnitude sample efficiency improvement compared to the already efficient Dreamer. We also refer the reader to Figure 3(b) (which is a clean illustration of the same data presented in Appendix F.6) for an example of the interpretable world model we can extract from MEAD.
For reference, we plot the performance for SOTA algorithms in the base MDP (without the Ab-MDP) as triangles and stars. In particular in MiniHack tasks (Fig.4(b)), model-free methods with random and global exploration bonuses (IMPALA, RND, ICM) perform poorly even after 50 million steps, and methods using episodic exploration bonus perform better (E3B). Nonetheless, all of them are less efficient than the least efficient method trained on the Ab-MDP.
These results show that Ab-MDP helps simplify the problem, with all methods reaching reasonable performances—a feat not achievable by most methods trained in the low level MiniHack MDP (Henaff et al.,, 2022). However, we also see that the problem is still non-trivial, with the model-free baseline still requiring millions of frames to achieve success. Our method reaches the same peak performance as the model-free method while using fewer samples than the model-based baseline.
@@ 4.2 Transfer and Compositions
Figure 5: Transfer experiments and a difficult compositional environment.
@@
We first place the agent in a “sandbox” environment where all objects are spawned with some probability (Appendix D.2.1). The agent is allowed to interact with the environment with only an exploration reward (no task-specific reward) for 200k frames. We then fine-tune the agent in each of the environments. For our model, we continue to train $f_{\theta}$ within the new environment. For Dreamer-v3, we transfer only the world model (encoder, decoder and dynamics function) and re-initialize the policy and reward functions. [^3] The point of this experiment is to evaluate world knowledge transfer. Nevertheless, fine-tuning the exploratory reward function and policy also achieves no better performance. Results for all five environments follow similar trends, thus we show a representative example in Figure 5(a) and present the full set in Figure 16(a).
Surprisingly, Dreamer-v3 trained with both count- and disagreement-based intrinsic rewards fails to get any meaningful rewards in the allocated fine-tuning budget (Figure 5(a), red & orange curves), although training for longer seems to show a some reward—indicating negative transfer (see Appendix F.1.2). Similarly, the model-free PPO policy shows negative transfer—reaching lower performance with the same data budget as compared to training from scratch (results in Appendix F.1.1 since x-axis in different order of magnitude). A possible explanation is while the agent observe all object types, the data distribution is dominated by seeing more objects on average in the training environment than in the evaluation environments, thus presenting a distribution shift. In contrast, our model exhibits good zero-shot performance, and reaches optimal performance efficiently.
To help the Dreamer-v3 baseline further, we give it a reward function to “use” the items seen in the pre-training sandbox environment. This is a kind of privileged information which biases the model representation to better model the evaluation-time task. We observe that this variant of Dreamer-v3 exhibits positive transfer, hinting at Dreamer’s representatin reliance on the reward function. Nevertheless, it does not exhibit the same degree of zero-shot or fine-tuning capability as our model.
@@
We further evaluate the ability of the models to transfer to new object types. For example, would a model learn to use a potion better, after it has learned to use a freeze wand, since the abstract level interactions of these objects are similar (the agent can stand on top of these objects, pick it up, and use them)? We take 200k frame checkpoints in the Freeze Random environment and re-train the model in the other 4 environments for an additional 50k frames.
@@
Finally, we designed an environment similar to the 5 difficult Minihack exploration levels, but with the required planning depth twice as long. [^5] Previous environments need a minimum of three abstract transitions to complete; this one needs at least six. This is an even more challenging task where the agent needs to pick up both a freeze and a levitate object, make itself levitate (using the levitate object), then cast a freezing spell to terminate the episode. Further, the levitate spell must be cast after all the objects have been picked up (since the agent cannot pickup objects while levitating), but before the freeze spell is cast. This is a challenging environment requiring precise sequences of abstract behaviours.
@@ 4.4.2 Count-based exploration
We observe using the count-based intrinsic reward (Equation 4) with MCTS was important for exploration. Figure 8 (right) shows count-based exploration discovering many more valid object attribute transitions compared to a randomly exploring agent. Correspondingly, this leads to a model $f_{\theta}$ that is better at achieving goals (Figure 8, left).
@@
Instead of fitting a discriminative model to success probability, we can directly use the empirical count-based estimate from a dataset (a non parametric model). This is done to learn the model in Zhang et al., (2018). We compare this choice in Appendix F.1.3 and show our parametric model has better sample efficiency, demonstrating the generalization benefits of our parametric forward model.
@@ 5 Related Work
Our Ab-MDP is a hierarchical state and temporal abstraction formalism and can be interpreted as Options (Appendix B.1). Within the Options Framework (Sutton et al.,, 1999; Precup,, 2000), Option discovery remains an open and important question. Existing methods typically do option discovery in a “bottom-up” fashion by learning structures from their training environments (Bacon et al.,, 2016; Machado et al.,, 2017; Ramesh et al.,, 2019; Machado et al.,, 2021). In contrast, our Ab-MDP uses “top-down” prior knowledge in the form of object-perception, and define options that perturb objects’ attributes.
@@
Abstraction is a fundamental concept (Abel,, 2022) and we select a subset of most relevant works. The Ab-MDP can be viewed as a more structured MDP formulation in which learning and planning is simplified. This relates to OO-MDP (Diuk et al.,, 2008; Keramati et al.,, 2018), which describes state as a set of object and object states which interact through relations. Also related is MaxQ and AMDP (Dietterich,, 1998; Gopalan et al.,, 2017), which provides a decomposition of a base MDP into multi-level sub-MDP hierarchies to plan more efficiently. Transition dynamics can also be defined as symbolic domain specific language (Mao et al.,, 2023). Above works require pre-specifying additional interaction rules, functions, and object relations to simplify planning, whereas we only assume the ability to see objects and their attributes, and learn the forward model as a neural network. More distantly related is Veerapaneni et al., (2019) which focus more on learning good object-centric representations rather than effective usage of an abstract level. Nasiriany et al., (2019) tackles sub-policy learning and planning with abstraction, focusing on implicitly representing abstractions rather than using semantically meaningful ones. The DAC-MDP (Shrestha et al.,, 2020) clusters low level states into “core states” to solve by value iteration, with a focus on getting good policies in an offline RL setting rather than abstract world model learning.
@@
Most similar is Zhang et al., (2018), which does model-based planning over abstract attributes. We differ in having a more expressive abstraction framework: modelling a flexible set of objects and shareable attributes, rather than a fixed, binary set of attributes only. We also use a parametric model to more efficiently learn and is able to transfer object information to new environments. We compare against their non-parametric approach and show better sample efficiency in Section 4.4.3.
@@
RL in semantic spaces As language models have grown in interest and sophistication, there has been an increasing interest in semantic and linguistic abstractions in RL. Early works such as Andreas et al., (2017) looked at handcrafted abstract spaces such as we do here to show that creating abstractions allows for better learning and exploration. Other works such as Jiang et al., (2019); Chen et al., (2021) specifically use natural language as the abstract space. In others, the language space abstraction is used as a way of guiding exploration Tam et al., (2022); Guo et al., (2023).
@@
Our work is related to other work in Model-based Reinforcement Learning (Sutton,, 1990; Kaiser et al.,, 2019; Doya et al.,, 2002). In particular we compare to the Dreamer line of work (Hafner et al.,, 2019, 2023; Sekar et al.,, 2020; Mendonca et al.,, 2021; Hafner et al.,, 2022). These methods were developed for a pixels setting and employ a two step learning strategy: learning a generative forward model to predicts the next states $X^{\prime}$, then learning a policy in imagination. We compare directly to this in our experiments, where we show that our method (which employs a discriminative forward model) learns more quickly and transfers better in our setting.
@@
Exploration is a well-studied problem in RL. For a more extensive literature review, Amin et al., (2021) gives a comprehensive picture. Most relevant to here are count-based methods such as Raileanu and Rocktäschel, (2020); Zhang et al., (2021). Our work takes inspiration from the hard exploration problems from  Henaff et al., (2022). There are also many disagreement-based methods including Sekar et al., (2020) and Mendonca et al., (2021); we compare directly to the former for our Dreamer benchmarks with pre-training.

### v3

@@ Abstract
We demonstrate the model’s ability to (i) efficiently solve single tasks, (ii) transfer zero-shot and few-shot across item types and environments, and (iii) plan across long horizons. Across a suite of 2D crafting and MiniHack environments, we empirically show our model significantly out-performs state-of-the-art low-level methods (without abstraction), as well as performant model-free and model-based methods using the same abstraction. Finally, we show how to learn low level object-perturbing policies via reinforcement learning, and the object mapping itself by supervised learning.
@@ 1 Introduction
Since the inception of reinforcement learning (RL), the problems of exploration and world model learning have been active, open questions. RL requires an agent to learn both basic motor abilities, and explore long sequences of interactions in the environment. Currently, we have made tremendous progress on problems of low-level control for short-horizon problems. With well defined single tasks, given enough samples (or demonstrations), there are well-developed ways of training low level policies reliably (Schulman et al., 2017; Hafner et al., 2023).
Perhaps a better way to explore complex environments is to treat the high-level exploration problem separately. We increasingly have the ability to get semantically rich abstractions: via representation learning (Locatello et al., 2020), object segmentation (Kirillov et al., 2023), and visual-language models  (Alayrac et al., 2022); even in difficult control settings such as robotics (Liu et al., 2024). In addition, with the recent explosion in the field of natural language processing, people are increasingly interested in addressing RL environments on a semantic level (Andreas et al., 2017; Jiang et al., 2019; Chen et al., 2021). Thus, we ask a timely question: if the agent has a good ability to perceive objects, what more can we do other than tabula rasa learning at the level of pixels and primitive motor actions?
@@
The main results of this work investigates decision making at the semantic level as defined by the Ab-MDP, assuming access to an object centric map and object perturbing policies. We empirically evaluate our method in a set of 2D crafting games, and MiniHack environments that are known to be very difficult to explore (Henaff et al., 2022). We then show how to learn components of the Ab-MDP when it is not given: by reinforcement learning object perturbation policies given an object mapping, and supervised learning of the object centric mapping itself. We leave the unsupervised discovery of the object centric mapping as complementary and future work.
@@ 2 Problem Setting
We begin by considering a (reward free) “base” Markov Decision Process (MDP) as the tuple $\langle\mathcal{S},\mathcal{A},P\rangle$, with state space $\mathcal{S}$, primitive action space $\mathcal{A}$, and transition probability function $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ (Puterman, 1994). Below we construct the “abstracted” space.
@@ 2.2 Abstracted Item-Attribute MDP
Specifically, the Ab-MDP is the tuple $\langle\mathcal{X},\mathcal{B},\textbf{T}\rangle$, with abstract states $\mathcal{X}$, abstract behaviours $\mathcal{B}$, and transition probability function $\textbf{T}:\mathcal{X}\times\mathcal{B}\times\mathcal{X}\rightarrow[0,1]$. $\textbf{T}(X^{\prime}|X,b)$ describes the probability of being in next abstract state $X^{\prime}$, when starting from $X$, executing behaviour $b$, and waiting until an abstract transition occurs. Figure 1 provides an example of an abstract transition and associated low level steps. The Ab-MDP can be interpreted via the Options Framework (Sutton et al., 1999; Precup, 2000), which we discuss in Appendix B.1.1.
@@ 3 Methods
While any standard RL methods can be used to solve the Ab-MDP, we develop a method to make better use of the item-attribute structure and item-perturbing behaviours. Specifically, we design a model-based method that efficiently explores and learn an accurate full world model. Our method, Model-based Exploration of abstracted Attribute Dynamics (MEAD) can be broken down into three components: (i) A forward model $f_{\theta}(X,b)$, (ii) A reward function $r(X,b)$, (iii) A planner which maximizes the reward function.
@@ 3.1 Forward model
A forward model takes in a current abstract state and behaviour to model the distribution over next states, $X_{t+1}\sim p(X_{t+1}|X_{t},b_{t})$. Typically, this distribution has support over the full state space $\mathcal{X}$. We observe that by construction, Ab-MDP considers competent policies as ones that change single item-attributes. Thus, we assume after an abstract transition, either (i) a single item’s attribute changes according to $b_{t}$, or (ii) the abstract state is unchanged. It is of course possible for multiple items’ attributes to change at once, or for an unexpected item’s attribute to change—we show with a competent set of policies and re-planning our model is robust to this in practice in Appendix B.2.2.
@@
We use model function class $f_{\theta}:\mathcal{X}\times\mathcal{B}\rightarrow[0,1]$ with trainable parameters $\theta$, and fit it to approximate success probability (Equation 1). We can estimate the empirical success probability from a dataset of trajectories simply by counting, and fit a binary cross entropy loss. We found this to be a stable, simple-to-optimize objective. More details in Appendix C.1.
@@ 3.2 Planning for exploration
(b) Dijkstra planning at eval time to maximize success probability of reaching any goal state
@@
We want the agent to explore indefinitely to discover how the world works and represent it in its model $f_{\theta}$. We use a count-based intrinsic reward introduced in Zhang et al. (2018) which encourages the agent to uniformly cover all state-behaviours (details in Appendix C.3):
@@ 4 Results
We use two sets of of environment. One is 2D crafting, which requires agent to craft objects following a Minecraft-like crafting tree with each crafting stage requiring multiple low level actions (Chen et al., 2020). Second is MiniHack (Samvelyan et al., 2021), which contains a series of extremely difficult exploration games (Henaff et al., 2022). We detail the environments further in Appendix D.
For baselines, we use strong model-free (asynch PPO, (Schulman et al., 2017; Petrenko et al., 2020)) and model-based (Dreamer-v3 (Hafner et al., 2023; Sekar et al., 2020) and MuZero (Schrittwieser et al., 2020; Werner Duvaud, 2019)) baselines. Dreamer-v3 learns a generative world model and plans using model-free method in imagination, while MuZero learns a generative world model and plans using MCTS. We outline different baseline methods in greater details and do additional comparisons in Appendix E. All the methods are trained in the Ab-MDP. We report low level steps and plot the 95% confidence interval. See Appendix G for other experimental details.
@@ 4.1 Learning from scratch in single environments
Figure 4: Results in games trained from scratch in Ab-MDP. Triangles and stars denote low-level only methods. For MiniHack, we also show the final performance of state-of-the-art exploration methods in the low level MDP (star/triangle), from Henaff et al. (2022). X-axis on log scale.
Figure 4 shows the from-scratch Ab-MDP training results for strong baselines and our method. We see that Dreamer-v3, a state-of-the-art model-based method, is highly sample efficient. MuZero, another model-based method utilizing online planning, is less efficient. PPO, a model-free method, stably optimizes the reward, but at a data budget of an additional $\sim$ two orders of magnitude (hundreds of thousands versus 10 million). Our method learns the Ab-MDP with an additional order of magnitude sample efficiency improvement compared to the already efficient Dreamer. We also refer the reader to Figure 3(b) (which is a clean illustration of the same data presented in Appendix F.6) for an example of the interpretable world model we can extract from MEAD.
For reference, we plot the performance for SOTA algorithms in the low level MDP (without Ab-MDP) as triangles and stars. In particular in MiniHack tasks (Fig.4(b)), model-free methods with random and global exploration bonuses (IMPALA, RND, ICM) perform poorly even after 50 million steps, and methods using episodic exploration bonus perform better (E3B). Nonetheless, all of them are less efficient than the least efficient method trained on the Ab-MDP.
These results show that Ab-MDP helps simplify the problem, with all methods reaching reasonable performances—a feat not achievable by most methods trained in the low level MiniHack MDP (Henaff et al., 2022). However, we also see that the problem is still non-trivial, with the model-free baseline still requiring millions of frames to achieve success. Our method reaches the same peak performance as the model-free method while using fewer samples than the model-based baseline.
@@ 4.2 Transfer and Compositions
Figure 5: Transfer experiments and a difficult compositional environment. Dotted lines in 5(a) and 5(b) are the same curves from Figure 4(b), with just the means shown for clarity.
@@
We first place the agent in a “sandbox” environment where all objects are spawned with some probability (Appendix D.2.1). The agent is allowed to interact with the environment with only an exploration reward (no task-specific reward) for 200k frames. We then fine-tune the agent in each of the environments. For our model, we continue to train $f_{\theta}$ within the new environment. For Dreamer-v3, we transfer only the world model (encoder, decoder and dynamics function) and re-initialize the policy and reward functions. [^3] The point of this experiment is to evaluate world knowledge transfer. Nevertheless, fine-tuning the exploratory reward function and policy also achieves no better performance. Results for all five environments follow similar trends; we show a representative example in Figure 5(a), with full results in Figure 16(a).
Surprisingly, Dreamer-v3 trained with both count- and disagreement-based intrinsic rewards fails to get any meaningful rewards in the allocated fine-tuning budget (Figure 5(a), red & orange curves), although training for longer seems to show a some reward—indicating negative transfer (see Appendix F.1.2). Similarly, the model-free PPO policy shows negative transfer—reaching lower performance with the same data budget as compared to training from scratch (results in Appendix F.1.1 since x-axis in different order of magnitude). A possible explanation is while the agent observes all object types, the data distribution is dominated by seeing more objects on average in the training environment than in the evaluation environments, thus presenting a distribution shift. In contrast, our model exhibits good zero-shot performance, and reaches optimal performance efficiently.
To help the Dreamer-v3 baseline further, we give it a reward function to “use” the items seen in the pre-training sandbox environment. This is a kind of privileged information which biases the model representation to better model the evaluation-time task. We observe that this variant of Dreamer-v3 exhibits positive transfer, hinting at Dreamer’s representation reliance on the reward function. Nevertheless, it does not exhibit the same degree of zero-shot or fine-tuning capability as our model.
@@
We further evaluate the ability of the models to transfer to new object types. For example, would a model learn to use a potion better, after it has learned to use a freeze wand, since the abstract level interactions of these objects are similar (the agent can stand on top of these objects, pick it up, and use them)? We take 200k frame checkpoints in the Freeze Random environment and train them in the other 4 environments for an additional 50k frames.
@@
Finally, we designed an environment similar to the 5 difficult Minihack exploration levels, but with the required planning depth twice as long. [^5] Previous environments need a minimum of three abstract transitions to complete; this one at least six. This is an even more challenging task where the agent needs to pick up both a freeze and a levitate object, make itself levitate (using the levitate object), then cast a freezing spell to terminate the episode. Further, the levitate spell must be cast after all the objects have been picked up (since the agent cannot pickup objects while levitating), but before the freeze spell is cast. This is a challenging environment requiring precise sequences of abstract behaviours.
@@ 4.4.2 Count-based exploration
We observe using the count-based intrinsic reward (Equation 4) with MCTS was important for exploration. Figure 8 shows count-based exploration discovering many more valid object attribute transitions compared to a randomly exploring agent. Correspondingly, this leads to a model $f_{\theta}$ that is better at achieving goals.
@@
Instead of fitting a discriminative model to success probability, we can directly use the empirical count-based estimate from a dataset (a non parametric model). This is done to learn the model in Zhang et al. (2018). We compare this choice in Appendix F.1.3 and show our parametric model has better sample efficiency, demonstrating the generalization benefits of our parametric forward model.
@@ 5 Related Work
Our Ab-MDP is a hierarchical state and temporal abstraction formalism and can be interpreted under the Options Framework (Appendix B.1). Within the Options Framework (Sutton et al., 1999; Precup, 2000), Option discovery remains an open and important question. Existing methods typically do option discovery in a “bottom-up” fashion by learning structures from their training environments (Bacon et al., 2016; Machado et al., 2017; Ramesh et al., 2019; Machado et al., 2021). In contrast, our Ab-MDP uses “top-down” prior knowledge in the form of object-perception, and define options that perturb objects’ attributes.
@@
Abstraction is a fundamental concept (Abel, 2022) and we select a subset of most relevant works. The Ab-MDP can be viewed as a more structured MDP formulation in which learning and planning is simplified. This relates to the Factored MDP (Boutilier et al., 1995; 1999; Guestrin et al., 2003; Degris & Sigaud, 2013), which models the conditional independence structure between factors in a factorized state space as Bayesian Networks. It also relates to OO-MDP (Diuk et al., 2008; Keramati et al., 2018), which describes state as a set of object and object states which interact through relations. We discuss their mathematical relationships in Appendix B.1.3. Also related is MaxQ and AMDP (Dietterich, 1998; Gopalan et al., 2017), which provides a decomposition of a base MDP into multi-level sub-MDP hierarchies to plan more efficiently. Transition dynamics can also be defined as symbolic domain specific language (Mao et al., 2023). Above works require pre-specifying additional interaction rules, functions, and object relations to simplify planning, whereas we only assume the ability to see objects and their attributes, and learn the forward model as a neural network. The CEE-US model of Sancaktar et al. (2022) use ensembles of graph neural networks to exploit an object-centric state representation which factorizes into entity and agent information, and show strong model learning and zero-shot generalization capabilities in a robot manipulation domain. DAC-MDP (Shrestha et al., 2020) clusters low level states into “core states” to solve by value iteration, with a focus on getting good policies in an offline RL setting rather than abstract world model learning. More distantly related is Veerapaneni et al. (2019) which focus more on learning good object-centric representations rather than effective usage of an abstract level. Nasiriany et al. (2019) tackles sub-policy learning and planning with abstraction, focusing on implicitly representing abstractions rather than using semantically meaningful ones.
Zhou et al. (2022); Haramati et al. (2024); Chang et al. (2023) are works that leverage object-centric representations with a set of discrete, factorized descriptions—a beneficial inductive bias Ab-MDP shares (i.e. item identities and attributes). In particular, Chang et al. (2023) uses slot attention to learn factorized “entity types” and “entity states”, albeit relying on more restrictive assumptions about entity dynamics. Haramati et al. (2024) extracts an object-centric representation using Deep Latent Particles, which extracts latent “particles” describing pixel-based objects in terms of their pixel location, size, etc. Both works offer potential ways of learning the object-centric map M. We discuss their relationship with Ab-MDP further in Appendix B.1.3 and compare against the model-learning method from Chang et al. (2023) in Appendix F.1.3.
@@
Most similar is Zhang et al. (2018), which does model-based planning over abstract attributes. We differ in having a more expressive abstraction framework: modelling a flexible set of objects and shareable attributes, rather than a fixed, binary set of attributes only. We also use a parametric model to more efficiently learn and is able to transfer object information to new environments. We compare against their non-parametric approach and show better sample efficiency in Section 4.4.3.
@@
RL in semantic spaces As language models have grown in interest and sophistication, there has been an increasing interest in semantic and linguistic abstractions in RL. Early works such as Andreas et al. (2017) looked at handcrafted abstract spaces such as we do here to show that creating abstractions allows for better learning and exploration. Other works such as Jiang et al. (2019); Chen et al. (2021) specifically use natural language as the abstract space. In others, the language space abstraction is used as a way of guiding exploration (Tam et al., 2022; Guo et al., 2023).
@@
Our work is related to other work in Model-based Reinforcement Learning (Sutton, 1990; Kaiser et al., 2019; Doya et al., 2002). In particular we compare to the Dreamer line of work (Hafner et al., 2019; 2023; Sekar et al., 2020; Mendonca et al., 2021; Hafner et al., 2022). These methods were developed for a pixels setting and employ a two step learning strategy: learning a generative forward model to predicts the next states $X^{\prime}$, then learning a policy in imagination. We compare directly to this in our experiments, where we show that our method (which employs a discriminative forward model) learns more quickly and transfers better in our setting.
@@
Exploration is a well-studied problem in RL. For a more extensive literature review, Amin et al. (2021) gives a comprehensive picture. Most relevant to here are count-based methods such as Raileanu & Rocktäschel (2020); Zhang et al. (2021). Our work takes inspiration from the hard exploration problems from  Henaff et al. (2022). There are also many disagreement-based methods including Sekar et al. (2020) and Mendonca et al. (2021); we compare directly to the former for our Dreamer benchmarks with pre-training.